# A rapid channel for the collisional formation and gravitational wave driven mergers of supermassive black hole seeds at high redshift

### v1

@@ Abstract
Motivated by JWST observations of dense, clumpy and clustered high redshift star formation, we simulate the hierarchical assembly of nine $M_{\mathrm{cl}}={10^{6}}\>\mathrm{M_{\odot}}$ star clusters using the BIFROST N-body code. Our low metallicity models ($0.01Z_{\odot}$) with post-Newtonian equations of motion for black holes include evolving populations of single, binary and triple stars. Massive stars grow by stellar collisions and collapse into intermediate mass black holes (IMBHs) up to $M_{\mathrm{\bullet}}\sim{6200}\>\mathrm{M_{\odot}}$, stellar multiplicity boosting the IMBH masses by a factor of 22223333. The IMBHs tidally disrupt (TDE) $\sim 50$ stars in $10$ Myr with peak TDE rates of $\Gamma\sim 10^{-5}$ yr-1 per cluster. These IMBHs are natural seeds for supermassive black holes (SMBHs) and the hierarchical assembly frequently leads to $>2$ SMBH seeds per cluster and their rapid mergers ($t<10$ Myr). We propose that a gravitational wave (GW) driven merger of IMBHs with ${1000}\>\mathrm{M_{\odot}}\lesssim M_{\bullet}\lesssim{10000}\>\mathrm{M_{\odot}}$ with comparable masses is a characteristic GW fingerprint of SMBH seed formation at redshifts $z>10$, and IMBH formation in similar environments at lower redshifts. Massive star clusters provide a unique environment for the early Universe GW studies for the next-generation GW observatories including the Einstein Telescope and the Laser Interferometer Space Antenna.
@@ 1 Introduction
[…]
@@
Out of the various proposed scenarios for the SMBH seed formation (Rees, 1984), the collisional runaway (e.g. Gold et al. 1965; Spitzer & Saslaw 1966) in dense and massive star clusters is an attractive scenario for the SMBH origins as it relies on relatively well understood physical processes and requires little fine tuning. Recent observations by the JWST have revealed dense, clumpy star formation as well as bound $\gtrsim{10^{6}}\>\mathrm{M_{\odot}}$ sub-pc size star clusters at $6\lesssim z\lesssim 10$ (e.g. Fujimoto et al. 2024; Adamo et al. 2024). Furthermore, observed high-z nitrogen enrichment at low metallicities and high star formation densities suggest the presence of dense environments in which stellar collisions can frequently occur (e.g. Topping et al. 2024; Isobe et al. 2025). This view of dense and clustered low metallicity star formation is further supported by Milky Way archaeology (Belokurov & Kravtsov, 2022). Moreover, high-resolution or star-by-star hydrodynamical simulations of low metallicity clouds (see e.g. Howard et al., 2018; Polak et al., 2024b) and galactic star-bursts (Lahén et al., 2020) indicate that massive star cluster formation is an inherently hierarchical process. Indeed, the regions of most intense star formation in the local Universe such as R136 in the Large Magellanic Cloud (Fahrion & De Marchi, 2024) and in the Antennae galaxies (Whitmore et al., 2010) host complex, hierarchical assemblies of star clusters, not isolated monolithically collapsed objects. Still, most numerical work on the stellar runaway scenario (e.g. Portegies Zwart et al. 1999, 2004; Gürkan et al. 2004; Freitag et al. 2006; Mapelli 2016; Sakurai et al. 2017; Rizzuto et al. 2021; Arca Sedda et al. 2023; González Prieto et al. 2024) has focused on isolated, spherical star cluster models or idealised few-cluster setups (Fujii & Portegies Zwart, 2014). Very recent models have introduced collisional stellar dynamics into star-by-star hydrodynamical simulations (e.g. Fujii et al. 2024; Polak et al. 2024a; Cournoyer-Cloutier et al. 2024; Lahén et al. 2025).
@@
In Rantala et al. (2024) we showed using direct N-body simulations that the hierarchical star cluster assembly with single stars boosts IMBH formation compared to monolithically formed clusters of the same mass and final structure. Massive stars can grow up to $\sim{2000}\>\mathrm{M_{\odot}}$ through stellar collisions, collapse into IMBHs and further grow by tidal disruption events (TDEs) and GW driven BH mergers. In this Letter, present a new suite of nine hierarchical N-body simulations of massive star cluster formation including initial (primordial) populations of binary and triple stars. Based on our simulation results we propose that a close to equal mass GW merger in the IMBH mass range is a characteristic GW fingerprint of the hierarchical runaway collisional scenario for the SMBH seed formation. Crucially, rapid IMBH-IMBH mergers frequently occur within $<10$ Myr of the cluster formation.
@@
This Letter is structured as follows. We present our N-body simulation code BIFROST and the numerical setup in section 2. In sections 3 and 4 we report our main stellar collision and GW merger results, and summarise in section 5.
@@ 2 Methods and initial conditions
For the hierarchical models we consider setups with three different stellar multiplicity properties (singles only, including binaries, including binaries and triples), generating three random realisations for each setup. The nine hierarchical cluster assembly regions consist of up to $900$ star clusters sampled from the universal slope $-2$ power law star cluster mass function (e.g. Elmegreen & Efremov 1996) up to maximum sub-cluster mass of $M_{\mathrm{cl}}={2.5\times 10^{5}}\>\mathrm{M_{\odot}}$. The most massive cluster is placed at the origin at zero velocity while the infalling clusters inhabit a spherical volume of $r=50$ pc in radius with $v_{\mathrm{r}}=-3.5$ km/s and a random velocity component of the same magnitude (see Rantala et al. 2024 for details). The total mass of each hierarchical model is $\sim{1.0\times 10^{6}}\>\mathrm{M_{\odot}}$ with $N\sim 1.8\times 10^{6}$ stars in the initial conditions. The hierarchical models are run until $t=10$ Myr after which most sub-clusters have merged into a single cluster typically containing multiple IMBHs in the mass range of $M_{\bullet}\sim{10^{3}}\>\mathrm{M_{\odot}}$${10^{4}}\>\mathrm{M_{\odot}}$.
@@ 3.1 Isolated cluster models
Figure 2: The maximum mass $m_{\mathrm{max}}$ of stellar collision products in isolated clusters with initial binaries (small crosses) and hierarchical (large symbols) setups as a function of star cluster mass $M_{\mathrm{cl}}$. The shaded area shows the isolated single star models of Rantala et al. (2024). With initial stellar binaries there is no suppression of collisions at $M_{\mathrm{cl}}\gtrsim 1.5\times{10^{5}}\>\mathrm{M_{\odot}}$. Stellar multiplicity in the models results in higher masses of stellar collision products by a factor of up to 2similar-toabsent2$\sim$ 2∼ 23333.
@@ 3.2 Hierarchically assembling cluster models
The three new hierarchical single star models are very similar to the hierarchical runs of Rantala et al. (2024). $3$-$7$ IMBHs form per assembled cluster, and the most massive IMBHs in each random realisation reach masses of $M_{\bullet}={1809}\>\mathrm{M_{\odot}}$, ${2708}\>\mathrm{M_{\odot}}$ and ${2671}\>\mathrm{M_{\odot}}$. Several IMBHs are ejected from the clusters via strong Newtonian few-body interactions with the other seeds, and at $t=10$ Myr two of the single star models still contain two or more IMBHs.
@@
Figure 5: The primary and secondary masses of BH binaries at $t=10$ Myr (left panel) and the GW mergers before $t<10$ Myr (right panel) in the hierarchical cluster assembly models. The Advanced LIGO / Advanced Virgo GW mergers in the GWTC-3 catalogue is shown as the cross symbols. Besides the GWTC-3 like stellar BH population, our hierarchical models result in IMBH-BH binaries and IMBH-BH GW mergers with ${300}\>\mathrm{M_{\odot}}\lesssim M_{\mathrm{\bullet,1}}\lesssim{3000}\>\mathrm{M_{\odot}}$ and $0.03\lesssim q\lesssim 0.1$. Most importantly, the hierarchical cluster assembly drives the formation of close to equal mass IMBH binaries and GW mergers with $M_{\mathrm{\bullet,1}}\gtrsim{2500}\>\mathrm{M_{\odot}}$ in less than $t=10$ Myr.
@@
For the more realistic hierarchical models including initial binaries, the merger trees of growing stars, IMBHs and their host clusters are illustrated in Fig. 3 for one of the models. Five sub-clusters form IMBHs through stellar collisions. In contrast to the single star setups, initial binary and triple populations allow for the formation in the most massive central star clusters as well. Three sub-clusters above the threshold mass of $M_{\mathrm{cl}}={10^{4}}\>\mathrm{M_{\odot}}$ above which collisions can frequently occur do not form BHs in the IMBH mass range. For all binary and triple models the first stellar collisions in the clusters occur early, before $t=0.5$ Myr, and almost always involve the members of a initial binary system, either colliding with each other or with other stars. By $t\sim 2.5$ Myr, the five sub-clusters have formed IMBHs in the range of $\sim{1000}\>\mathrm{M_{\odot}}\lesssim M_{\bullet}\lesssim{2550}\>\mathrm{M_{\odot}}$.
@@
After their formation, the IMBHs may subsequently grow by tidally disrupting (mostly) main sequence stars as well as evolved stars. Tidal disruption events can contribute up to $10\%$$15\%$ of the IMBH mass with most of the TDEs occurring soon after the formation of the IMBHs. The number of TDEs by IMBHs per assembled cluster (40similar-toabsent40$\sim$ 40∼ 4050505050 for $M_{\bullet}\gtrsim{100}\>\mathrm{M_{\odot}}$ and (30similar-toabsent30$\sim$ 30∼ 3045454545 for $M_{\bullet}\gtrsim{1000}\>\mathrm{M_{\odot}}$) in our models show no strong dependence on initial stellar multiplicity. The rate of TDEs by $M_{\bullet}\geq{100}\>\mathrm{M_{\odot}}$ IMBHs peaks at $\Gamma\sim 2\times 10^{-5}$ yr-1 per assembling cluster at $2.5$ Myr $\lesssim t\lesssim 3.75$ Myr. At later times, the central density of the formed cluster decreases due to mass loss and dynamical relaxation and the TDE rate accordingly declines by a factor of 555510101010. During the $10$ Myr of the simulations, IMBH growth by GW mergers with stellar mass BHs is less important with only few ($<5$) IMBH-BH merger occurring in total in the models of this study.
@@
IMBHs formed in the infalling sub-clusters rapidly sink to the main assembling cluster by the hierarchical star cluster merger process in $\lesssim 10$ Myr, of the order of $\sim 1$ dynamical time $t_{\mathrm{dyn}}\sim 5.3$ Myr of the initial system (see also Shi et al. 2024). In total 5 IMBHs formed through stellar collisions end up in the assembled cluster or its outskirts before $t=10$ Myr. A hierarchical triple IMBH interaction ($M_{\bullet}={2910}\>\mathrm{M_{\odot}},{2682}\>\mathrm{M_{\odot}}$ and ${1968}\>\mathrm{M_{\odot}}$) at 55556666 Myr drives the most massive IMBH binary to a GW merger, the resulting BH with $M_{\bullet}\sim{5300}\>\mathrm{M_{\odot}}$ being ejected from the cluster with a GW recoil kick of $v_{\mathrm{kick}}=310$ km/s. IMBH-IMBH mergers may provide an important channel for the early growth of SMBH seeds without any AGN or TDE signature, but only if the seeds can be retained in their environments. Around $t=6$ Myr, two star clusters deliver the subsequent IMBHs of $M_{\bullet}={2479}\>\mathrm{M_{\odot}}$ and $M_{\bullet}={991}\>\mathrm{M_{\odot}}$ into the central cluster. By the end of the simulation at $t=10$ Myr, the more massive of these has sunk to the centre of the main cluster, forming a binary system with the pre-existing central IMBH ($M_{\bullet}={1985}\>\mathrm{M_{\odot}}$). The future GW merger and GW recoil of these two IMBHs determines whether the cluster will eventually contain an IMBH of $M_{\bullet}\sim{4000}\>\mathrm{M_{\odot}}$, $\sim{1000}\>\mathrm{M_{\odot}}$, or no IMBHs at all.
@@
Overall, the models with initial binaries form 55557777 IMBHs per assembled cluster with on average 22223333 times higher masses compared to the single star models. In the second hierarchical model with initial binaries several IMBHs form and one IMBH with $M_{\bullet}={542}\>\mathrm{M_{\odot}}$ is ejected from the final cluster in a strong Newtonian few-body interaction. The central IMBH binary at $t=10$ Myr has a mass ratio of $q\sim 1/4$ with IMBH masses of $M_{\bullet}={935}\>\mathrm{M_{\odot}}$ and $M_{\bullet}={3904}\>\mathrm{M_{\odot}}$. Another IMBH with $M_{\bullet}={638}\>\mathrm{M_{\odot}}$ from a cluster merger is slowly sinking at the outskirts of the cluster. The third hierarchical model with binaries produces the most massive GW merger of this study with close to equal IMBH masses of $M_{\bullet}={3516}\>\mathrm{M_{\odot}}$ and $M_{\bullet}={2980}\>\mathrm{M_{\odot}}$. The remnant IMBH with $M_{\bullet}={6190}\>\mathrm{M_{\odot}}$ became marginally unbound of the host cluster by the relativistic GW recoil kick of $v_{\mathrm{kick}}=107$ km/s. At $t=10$ Myr, three less massive IMBHs (${314}\>\mathrm{M_{\odot}}\lesssim M_{\bullet}\lesssim{545}\>\mathrm{M_{\odot}}$) sink in the cluster outskirts, originate from three individual cluster mergers, but have not yet had a sufficient time to reach the cluster centre.
@@
Finally, we focus on hierarchical clusters with initial triples. The total number of any stellar collisions is the highest in the triples models, and the triple models give rise to $9$-$11$ IMBHs per assembled cluster. Still, including initial triples in the models only moderately increases the final collision product masses, the most massive IMBH in each triple simulations reaching $M_{\bullet}={3374}\>\mathrm{M_{\odot}}$, ${4676}\>\mathrm{M_{\odot}}$ and ${2479}\>\mathrm{M_{\odot}}$. $\sim 30\%$ of the IMBHs are ejected before $t=10$ Myr via Newtonian interactions and GW recoil kicks. Initial stellar multiplicity results in earlier first collisions and a head start for the collision cascades compared to the single star models, leading to by a factor of 22223333 higher IMBH masses compared to models with no initial multiplicity. Further increased stellar multiplicity (triples vs binaries) results in a higher number of formed IMBHs per final cluster, but only mildly increases the IMBH masses.

### v2

@@ Abstract
Motivated by JWST observations of dense, clumpy and clustered high redshift star formation, we simulate the hierarchical assembly of nine $M_{\mathrm{cl}}={10^{6}}\>\mathrm{M_{\odot}}$ star clusters using the BIFROST N-body code. Our low metallicity models ($0.01Z_{\odot}$) with post-Newtonian equations of motion for black holes include evolving populations of single, binary and triple stars. Massive stars grow by stellar collisions and collapse into intermediate mass black holes (IMBHs) up to $M_{\mathrm{\bullet}}\sim{6200}\>\mathrm{M_{\odot}}$, stellar multiplicity boosting the IMBH masses by a factor of 2233. The IMBHs tidally disrupt (TDE) $\sim 50$ stars in $10$ Myr with peak TDE rates up to $\Gamma\sim 5\times 10^{-5}$ yr-1 per cluster. These IMBHs are natural seeds for supermassive black holes (SMBHs) and the hierarchical assembly frequently leads to >2>2 SMBH seeds per cluster and their rapid mergers ($t<10$ Myr). We propose that a gravitational wave (GW) driven merger of IMBHs with ${1000}\>\mathrm{M_{\odot}}\lesssim M_{\bullet}\lesssim{10000}\>\mathrm{M_{\odot}}$ with comparable masses is a characteristic GW fingerprint of SMBH seed formation at redshifts $z>10$, and IMBH formation in similar environments at lower redshifts. Massive star clusters provide a unique environment for the early Universe GW studies for the next-generation GW observatories including the Einstein Telescope and the Laser Interferometer Space Antenna.
@@ 1 Introduction
Gravitational wave (GW) driven black hole (BH) mergers in the intermediate-mass black hole (IMBH; Mezcua 2017; Greene et al. 2020) mass range of $M_{\bullet}\sim{10^{2}}\>\mathrm{M_{\odot}}$${10^{5}}\>\mathrm{M_{\odot}}$ at redshifts $z>6$ will be observable with the next-generation GW observatories such as the Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al. 2017), the Einstein Telescope (Punturo et al. 2010) and the Cosmic Explorer (Reitze et al. 2019). This high-z GW window provides an unprecedented opportunity to constrain the theoretical models for the origins of supermassive black holes (SMBHs), a major open question in modern astrophysics only deepened by the James Webb Space Telescope (JWST) observations of accreting SMBHs in active galactic nuclei (AGN) at $z>10$ (Maiolino et al., 2024). To explain the z=6z=677 AGN population (Fan et al., 2023; Inayoshi et al., 2020), $M_{\bullet}={10^{4}}\>\mathrm{M_{\odot}}$ SMBH seeds growing at the Eddington rate must have formed around $z\sim 10$$11$ to reach $M_{\bullet}={10^{9}}\>\mathrm{M_{\odot}}$ by $z=6$. By similar argument the progenitor of the GN-z11 SMBH ($M_{\bullet}\sim{1.6\times 10^{6}}\>\mathrm{M_{\odot}}$; Maiolino et al. 2024) should have formed at $z\sim 18$1919 if it initially had $M_{\bullet}={10^{4}}\>\mathrm{M_{\odot}}$.
@@
Out of the various proposed scenarios for the SMBH seed formation (Rees, 1984), the collisional runaway (e.g. Gold et al. 1965; Spitzer & Saslaw 1966) in dense and massive star clusters is an attractive scenario for the SMBH origins as it relies on relatively well understood physical processes and requires little fine tuning. Recent observations by the JWST have revealed dense, clumpy star formation as well as bound $\gtrsim{10^{6}}\>\mathrm{M_{\odot}}$ sub-pc size star clusters at $6\lesssim z\lesssim 10$ (e.g. Fujimoto et al. 2024; Adamo et al. 2024). Furthermore, observed high-z nitrogen enrichment at low metallicities and high star formation densities suggest the presence of dense environments in which stellar collisions can frequently occur (e.g. Topping et al. 2024; Isobe et al. 2025). This view of dense and clustered low metallicity star formation is further supported by Milky Way archaeology (Belokurov & Kravtsov, 2022). Moreover, high-resolution or star-by-star hydrodynamical simulations of low metallicity clouds (see e.g. Howard et al., 2018; Polak et al., 2024b) and galactic star-bursts (Lahén et al., 2020) indicate that massive star cluster formation is an inherently hierarchical process. Indeed, the regions of most intense star formation in the local Universe such as R136 in the Large Magellanic Cloud (Fahrion & De Marchi, 2024) and in the Antennae galaxies (Whitmore et al., 2010) host complex, hierarchical assemblies of star clusters, not isolated monolithically collapsed objects. Still, most numerical work on the stellar runaway scenario (e.g. Portegies Zwart et al. 1999, 2004; Gürkan et al. 2004; Freitag et al. 2006; Mapelli 2016; Sakurai et al. 2017; Rizzuto et al. 2021; Arca Sedda et al. 2023; González Prieto et al. 2024; Vergara et al. 2025) has focused on isolated, spherical star cluster models or idealised few-cluster setups (Fujii & Portegies Zwart, 2014). Very recent models have introduced collisional stellar dynamics into star-by-star hydrodynamical simulations (e.g. Fujii et al. 2024; Polak et al. 2024a; Cournoyer-Cloutier et al. 2024; Lahén et al. 2025).
@@
In Rantala et al. (2024) we showed using direct N-body simulations that the hierarchical star cluster assembly with single stars boosts IMBH formation compared to monolithically formed clusters of the same mass and final structure. Massive stars can grow up to $\sim{2000}\>\mathrm{M_{\odot}}$ through stellar collisions, collapse into IMBHs and further grow by tidal disruption events (TDEs) and GW driven BH mergers. In this Letter, present a new suite of nine hierarchical N-body simulations of massive star cluster formation including initial (primordial) populations of binary and triple stars. Based on our simulation results we propose that a close to equal mass GW merger in the IMBH mass range is a characteristic GW fingerprint of the hierarchical runaway collisional scenario for the SMBH seed formation. Crucially, rapid IMBH-IMBH mergers frequently occur within <10<10 Myr of the cluster formation.
@@
This Letter is structured as follows. We present our N-body simulation code BIFROST and the numerical setup in section 2. In sections 3 and 4 we report our main results, focusing on stellar collisions and GW mergers. Finally, we summarise in section 5. The structure, kinematics and the stellar multiplicity content of the assembled star clusters is analysed in detail in a separate study (Rantala et al., 2025).
@@ 2 Methods and initial conditions
For the hierarchical models we consider setups with three different stellar multiplicity properties (singles only, including binaries, including binaries and triples), generating three random realisations for each setup. The nine hierarchical cluster assembly regions consist of up to $900$ star clusters sampled from the universal slope 2-2 power law star cluster mass function (e.g. Elmegreen & Efremov 1996) up to maximum sub-cluster mass of $M_{\mathrm{cl}}={2.5\times 10^{5}}\>\mathrm{M_{\odot}}$. The most massive cluster is placed at the origin at zero velocity while the infalling clusters inhabit a spherical volume of $r=50$ pc in radius with $v_{\mathrm{r}}=-3.5$ km/s and a random velocity component of the same magnitude (see Rantala et al. 2024 for details). The total mass of each hierarchical model is $\sim{1.0\times 10^{6}}\>\mathrm{M_{\odot}}$ with $N\sim 1.8\times 10^{6}$ stars in the initial conditions. The hierarchical models are run until $t=10$ Myr after which most sub-clusters have merged into a single cluster typically containing multiple IMBHs in the mass range of $M_{\bullet}\sim{10^{3}}\>\mathrm{M_{\odot}}$${10^{4}}\>\mathrm{M_{\odot}}$.
@@ 3.1 Isolated cluster models
Figure 2: The maximum mass $m_{\mathrm{max}}$ of stellar collision products in isolated clusters with initial binaries (small crosses) and hierarchical (large symbols) setups as a function of star cluster mass $M_{\mathrm{cl}}$. The shaded area shows the isolated single star models of Rantala et al. (2024). With initial stellar binaries there is no suppression of collisions at $M_{\mathrm{cl}}\gtrsim 1.5\times{10^{5}}\>\mathrm{M_{\odot}}$. Stellar multiplicity in the models results in higher masses of stellar collision products by a factor of up to $\sim 2$33.
@@ 3.2 Hierarchically assembling cluster models
The three new hierarchical single star models are very similar to the hierarchical runs of Rantala et al. (2024). 33-77 IMBHs form per assembled cluster, and the most massive IMBHs in each random realisation reach masses of $M_{\bullet}={1809}\>\mathrm{M_{\odot}}$, ${2708}\>\mathrm{M_{\odot}}$ and ${2671}\>\mathrm{M_{\odot}}$. Several IMBHs are ejected from the clusters via strong Newtonian few-body interactions with the other seeds, and at $t=10$ Myr two of the single star models still contain two or more IMBHs.
@@
Figure 5: The primary and secondary masses of BH binaries at $t=10$ Myr (left panel) and the GW mergers before $t<10$ Myr (right panel) in the hierarchical cluster assembly models. The Advanced LIGO / Advanced Virgo GW mergers in the GWTC-3 catalogue is shown as the cross symbols. Besides the GWTC-3 like stellar BH population, our hierarchical models result in IMBH-BH binaries and IMBH-BH GW mergers with ${300}\>\mathrm{M_{\odot}}\lesssim M_{\mathrm{\bullet,1}}\lesssim{3000}\>\mathrm{M_{\odot}}$ and $0.03\lesssim q\lesssim 0.1$. Most importantly, the hierarchical cluster assembly drives the formation of close to equal mass IMBH binaries and GW mergers with $M_{\mathrm{\bullet,1}}\gtrsim{2500}\>\mathrm{M_{\odot}}$ in less than $t=10$ Myr. The dashed lines indicate logarithmically spaced fixed values for the total binary mass $M_{\mathrm{\bullet,1}}+M_{\mathrm{\bullet,2}}$.
@@
For the more realistic hierarchical models including initial binaries, the merger trees of growing stars, IMBHs and their host clusters are illustrated in Fig. 3 for one of the models. Five sub-clusters form IMBHs above $M_{\bullet}\gtrsim{1000}\>\mathrm{M_{\odot}}$ through stellar collisions. In contrast to the single star setups, initial binary and triple populations allow for the formation in the most massive central star clusters as well. Three sub-clusters above the threshold mass of $M_{\mathrm{cl}}={10^{4}}\>\mathrm{M_{\odot}}$ above which collisions can frequently occur do not form BHs in the IMBH mass range. For all binary and triple models the first stellar collisions in the clusters occur early, before $t=0.5$ Myr, and almost always involve the members of a initial binary system, either colliding with each other or with other stars. By $t\sim 2.5$ Myr, the five sub-clusters have formed IMBHs in the range of $\sim{1000}\>\mathrm{M_{\odot}}\lesssim M_{\bullet}\lesssim{2550}\>\mathrm{M_{\odot}}$.
@@
After their formation, the IMBHs may subsequently grow by tidally disrupting (mostly) main sequence stars as well as evolved stars. Tidal disruption events can contribute up to $10\%$$15\%$ of the IMBH mass with most of the TDEs occurring soon after the formation of the IMBHs. The number of TDEs by IMBHs per assembled cluster ($\sim 40$$50$ for $M_{\bullet}\gtrsim{100}\>\mathrm{M_{\odot}}$ and ($\sim 30$$45$ for $M_{\bullet}\gtrsim{1000}\>\mathrm{M_{\odot}}$) in our models show no strong dependence on initial stellar multiplicity. The rate of TDEs by $M_{\bullet}\geq{100}\>\mathrm{M_{\odot}}$ IMBHs peaks at $\Gamma\sim 4$$5\times 10^{-5}$ yr-1 per assembling cluster at $2.5$ Myr $\lesssim t\lesssim 3.75$ Myr. At later times, the central density of the formed cluster decreases due to mass loss and dynamical relaxation and the TDE rate accordingly declines by a factor of 551010 (see also Rizzuto et al. 2023). During the $10$ Myr of the simulations, IMBH growth by GW mergers with stellar mass BHs is less important with only few (<5<5) IMBH-BH merger occurring in total in the models of this study.
@@
IMBHs formed in the infalling sub-clusters rapidly sink to the main assembling cluster by the hierarchical star cluster merger process in $\lesssim 10$ Myr, of the order of $\sim 1$ dynamical time $t_{\mathrm{dyn}}\sim 5.3$ Myr of the initial system (see also Shi et al. 2024). In total 5 IMBHs formed through stellar collisions end up in the assembled cluster or its outskirts before $t=10$ Myr. A hierarchical triple IMBH (see also Liu et al. 2024) interaction ($M_{\bullet}={2910}\>\mathrm{M_{\odot}},{2682}\>\mathrm{M_{\odot}}$ and ${1968}\>\mathrm{M_{\odot}}$) at 5566 Myr drives the most massive IMBH binary to a GW merger, the resulting BH with $M_{\bullet}\sim{5300}\>\mathrm{M_{\odot}}$ being ejected from the cluster with a GW recoil kick of $v_{\mathrm{kick}}=310$ km/s. IMBH-IMBH mergers may provide an important channel for the early growth of SMBH seeds without any AGN or TDE signature, but only if the seeds can be retained in their environments. Around $t=6$ Myr, two star clusters deliver the subsequent IMBHs of $M_{\bullet}={2479}\>\mathrm{M_{\odot}}$ and $M_{\bullet}={991}\>\mathrm{M_{\odot}}$ into the central cluster. By the end of the simulation at $t=10$ Myr, the more massive of these has sunk to the centre of the main cluster, forming a binary system with the pre-existing central IMBH ($M_{\bullet}={1985}\>\mathrm{M_{\odot}}$). The future GW merger and GW recoil of these two IMBHs determines whether the cluster will eventually contain an IMBH of $M_{\bullet}\sim{4000}\>\mathrm{M_{\odot}}$, $\sim{1000}\>\mathrm{M_{\odot}}$, or no IMBHs at all.
@@
Overall, the models with initial binaries form 4466 IMBHs per assembled cluster with on average 2233 times higher masses compared to the single star models. In the second hierarchical model with initial binaries several IMBHs form and one IMBH with $M_{\bullet}={542}\>\mathrm{M_{\odot}}$ is ejected from the final cluster in a strong Newtonian few-body interaction. The central IMBH binary at $t=10$ Myr has a mass ratio of $q\sim 1/4$ with IMBH masses of $M_{\bullet}={935}\>\mathrm{M_{\odot}}$ and $M_{\bullet}={3904}\>\mathrm{M_{\odot}}$. Another IMBH with $M_{\bullet}={638}\>\mathrm{M_{\odot}}$ from a cluster merger is slowly sinking at the outskirts of the cluster. The third hierarchical model with binaries produces the most massive GW merger of this study with close to equal IMBH masses of $M_{\bullet}={3516}\>\mathrm{M_{\odot}}$ and $M_{\bullet}={2980}\>\mathrm{M_{\odot}}$. The remnant IMBH with $M_{\bullet}={6190}\>\mathrm{M_{\odot}}$ became marginally unbound of the host cluster by the relativistic GW recoil kick of $v_{\mathrm{kick}}=107$ km/s. At $t=10$ Myr, three less massive IMBHs (${314}\>\mathrm{M_{\odot}}\lesssim M_{\bullet}\lesssim{545}\>\mathrm{M_{\odot}}$) sink in the cluster outskirts, originating from three individual cluster mergers, but have not yet had a sufficient time to reach the cluster centre.
@@
Finally, we focus on hierarchical clusters with initial triples. The total number of any stellar collisions is the highest in the triples models, and the triple models give rise to 99-$11$ IMBHs per assembled cluster. Still, including initial triples in the models only moderately increases the final collision product masses, the most massive IMBH in each triple simulations reaching $M_{\bullet}={3374}\>\mathrm{M_{\odot}}$, ${4676}\>\mathrm{M_{\odot}}$ and ${2479}\>\mathrm{M_{\odot}}$. $\sim 30\%$ of the IMBHs are ejected before $t=10$ Myr via Newtonian interactions and GW recoil kicks. Initial stellar multiplicity results in earlier first collisions and a head start for the collision cascades compared to the single star models, leading to by a factor of 2233 higher IMBH masses compared to models with no initial multiplicity. Further increased stellar multiplicity (triples vs binaries) results in a higher number of formed IMBHs per final cluster, but only mildly increases the IMBH masses.